A factory costs $970,000. You reckon that it will produce an inflow after operating costs of $187,000 a year for 15 years.

a. If the opportunity cost of capital is 11%, what is the net present value of the factory?

b. What will the factory be worth at the end of eight years?
Answer #1

a)NPV

CF NPV = — - Initial Investment (1+rn

187,000 NPV = - 0.11 * (1 + 0.11 15 -- 970,000

NPV = 1,700,000 * 0.7909956533-970,000

NPV = $374,692.61061

b) At the end of eight years, the factory's worth is equal to the present value of future cash flows from year 9 to 15. A total of seven cash flows.

PV_{8} = \frac{CF_{9}}{r} * \left [ 1 - \frac{1}{(1+r)^{n}} \right ]

PV% =- 187, 000 0.11 * ΤΙ – | (1 + 0.11)

PV = 1,700,000 *0.5183415891

PV = $881, 180.70147

This is the worth of the factory at the end of 8 years.

Can you please upvote? Thank You :-)